OpenStudy (anonymous):

solve initial value problem and find interval of validity to the solution yprime=x^2+3x+2/y-2 y(1)=4

OpenStudy (anonymous):

\[y \prime = \frac{ x^2+3x+2 }{ y-2 } \] \[y(1)=4\]

OpenStudy (anonymous):

\[y \prime (y-2) = x^2+3x+2\]

OpenStudy (anonymous):

and I go to integrate both sides but I'm having a brain fart. would it be; \[\frac{ y^2 }{ 2 }-2y+x=\frac{ x^3 }{ 3 }+\frac{ 3x^2 }{ 2 }+2x\]

OpenStudy (anonymous):

+c* not x

OpenStudy (anonymous):

SCRATCH finding interval of validity, read wrong question